# Diff-SSL-G-COMP: Towards a large-scale and diverse dataset for virtual analog modeling

### v1

@@ Abstract
Virtual Analog (VA) modeling aims to simulate the behavior of hardware circuits via algorithms to replicate their tone digitally. Dynamic Range Compressor (DRC) is an audio processing module that controls the dynamics of a track by reducing and amplifying the volumes of loud and quiet sounds, which is essential in music production. In recent years, neural-network-based VA modeling has shown great potential in producing high-fidelity models. However, due to the lack of data quantity and diversity, their generalization ability in different parameter settings and input sounds is still limited. To tackle this problem, we present Diff-SSL-G-Comp, the first large-scale and diverse dataset for modeling the SSL 500 G-Bus Compressor. Specifically, we manually collected 175 unmastered songs from the Cambridge Multitrack Library. We recorded the compressed audio in 220 parameter combinations, resulting in an extensive 2528-hour dataset with diverse genres, instruments, tempos, and keys. Moreover, to facilitate the use of our proposed dataset, we conducted benchmark experiments in various open-sourced black-box and grey-box models, as well as white-box plugins. We also conducted ablation studies in different data subsets to illustrate the effectiveness of improved data diversity and quantity. The dataset and demos are on our project page: http://www.yichenggu.com/DiffSSLGComp/.
@@ 1 Introduction
Virtual Analog (VA) modeling aims to simulate analog audio devices digitally. Dynamic Range Compressor (DRC) is an audio processing module that compresses the dynamics of a track by reducing and amplifying the volumes of loud and quiet sounds, which is essential in music production [1]. VA modeling on DRC is important but always considered to be challenging due to its characteristics: non-linear, time-invariant, and long temporal dependency.
To model an analog compressor, early DSP-based methods utilized white-box models [1]. Although they can achieve outstanding performances, their model is usually device-specific and involves extensive human labor. As deep learning develops, neural-network-based black-box models [2, 3, 4, 5] have become popular in recent years due to their superior ability to model analog devices in a data-driven way, thus avoiding the high cost in human labor. More recently, following [6], grey-box models [7, 8] are also proposed, which combine the explainability and data-driven training from white and black-box models and achieve excellent performances.
Although these existing models have shown their potential in producing high-fidelity models, their generalization ability is still limited due to a lack of data quantity and diversity. In particular, existing VA modeling datasets [8, 9, 10, 11] primarily utilize synthetic test signals as inputs and only record scattered parameter combinations, which may significantly constrain the model performance when encountering real-world recordings and unseen parameters.
@@
To tackle this issue, this work presents Diff-SSL-G-Comp, the first large-scale and diverse dataset for modeling the SSL 500 G-Bus Compressor  [^1] https://solidstatelogic.com/products/stereo-bus-compressor-module. Specifically, we manually selected 175 unmastered real-world songs from the Cambridge Multitrack Library  [^2] https://www.cambridge-mt.com/ms/mtk/ and recorded the compressed signals in 220 parameter combinations, which results in an extensive 2528-hour dataset with diverse genres, instruments, tempos, and keys. To facilitate the use of our dataset, we conducted benchmarking experiments on various open-sourced black-box and grey-box models, as well as available white-box plugins. We also conducted ablation studies on data subsets with different songs and data scales to illustrate the effectiveness of improved data diversity and quantity.
2 Related Work
@@
Table 1: A comparison of Diff-SSL-G-Comp with existing VA modeling datasets regarding DRC.
@@
DSP-based white-box models generally comprise a gain computer and a level detector with different detailed algorithm designs [1, 12], which has been well-studied over the years. Apart from this, works have also been proposed to increase computational efficiency; in particular,  [13] utilized block processing strategies for speeding up computation, and  [14] implemented the module via FFTs to use its parallel computing ability. Automatic DRC parameter control is another relevant research topic. Specifically,  [15] investigated automatic parameter control based on different extracted features, followed by  [16] to further extend to multi-track scenarios. With the development of machine learning techniques,  [17] proposed a regression model to control the parameters based on a referenced sound, followed by  [18] using a more potent random forest algorithm. Based on these white-box models, various studies have been proposed to simulate analog compressors, including creating virtual models for specific devices [19] or detailed components like optocouplers [20] and amplifiers [21]. In general, white-box models have high explainability in model design and good computational efficiency. However, due to the non-linear distortion in analog circuits, extensive human expert labor is usually needed to obtain a high-fidelity model in a specific device with long-time tuning. This makes it very expensive and impractical compared with the data-driven black-box and grey-box methods.
neural-network-based black-box models have been developed a lot in recent years.  [10] first proposed an autoencoder model to model various audio effects.  [2] utilized the LSTM model for optimizing the long-term dependencies, followed by [22] to further expand into the hyper RNN model with an in-detailed comparison between RNN, GRU, and LSTM models. To utilize the advantages of CNNs,  [23] first employed the WaveNet [24] structure on digital audio effects. Based on this work,  [3] proposed a temporal convolutional network (TCN) with larger receptive fields and huge dilation factors.  [4] further improved this architecture by integrating the feature-wise linear modulation (FiLM) [25] layers in modeling the parameter conditions. State Space Model (SSM) [26] is another technique to model long-term dependent time series via decomposing a dynamic system into structured state variables.  [5] first employed the S4 blocks in VA modeling, obtaining outstanding performances, followed by  [27] further adopting the latest S6 model [28]. With the development of differentiable digital signal processing (DDSP) [6], works are also proposed to utilize the DSP models’ explainability and efficiency. Specifically,  [29] proposed the differentiable biquad filters for machine learning applications, and  [30] integrated the biquad filter modeling with Koopman Networks [31] to operate in a higher dimensional state space. These advances have also made the neural grey-box models viable. In particular,  [7] utilized the classic white-box DRC [1] design with MLPs predicting the parameters in each time frame, followed by  [32] to further simplify the model into parametric Gains for compression and supplementary EQs for non-linear distortion.
Despite the rapid development of VA models, the publicly available datasets are still scarce, with limited data quantity and diversity. Table 1 illustrates the details of the existing datasets regarding DRC. Specifically, early attempts [9] primarily processed short instrument recordings in a specific parameter setting for nonparametric models. SignalTrain [10] presented the first parametric dataset in modeling the optical compressor LA-2A. It used various randomly generated test signals and a few instrument recordings as the input signals and recorded 20 equally sampled parameter combinations. After that,  [11] proposed the CL-1B dataset with real-world recordings as inputs and more parameter combinations. Recent works like [27] also presented datasets with more diverse devices but often with limited data scale and parameter combinations.
Data scaling has been shown effective in many areas [33, 34]. For instance, Mert [35] utilized a music mixture of 160K hours to scale up a self-supervised representation learning model with 330M parameters, obtaining outstanding performance in music information retrieval; Yue [36] constructed a 650K hours music mixture to train a 7B parameter model for music generation, obtaining state-of-the-art (SOTA) performance; Stable Audio [37] collected 73k hours of audio recordings, leading to SOTA audio generation model with 1B parameters; Emilia [38] presented a 101K hours open-sourced speech dataset, facilitating SOTA speech generations models [39, 40, 41]. Following these previous works, this paper presents Diff-SSL-G-Comp, the first large-scale and diverse dataset in VA modeling. We aim to address the limitations in existing datasets and explore the effectiveness of the improved data scale.
3 Diff-SSL-G-Comp
@@
Figure 1: Duration statistics (hours) of the unmastered songs used as input signals in Diff-SSL-G-Comp by genres and instruments.
As discussed in Section 2, existing VA modeling datasets are undiversified with limited data scales, which may restrict the model performance. To address this issue, we present Diff-SSL-G-Comp, an extensive and diverse dataset for modeling the SSL 500 G-BUS Compressor \@footnotemark. This section provides the construction details, statistics, and analysis of Diff-SSL-G-Comp.
@@ 3.1 Dataset Construction
Diff-SSL-G-Comp comprises unmastered songs with different genres, instruments, tempos, and keys processed with varying compression parameters. In particular, we manually selected 175 unmastered songs from the Cambridge Multitrack Library \@footnotemark. We used Reaper  [^3] https://www.reaper.fm/ as the Digital Audio Workstation (DAW) to process the data automatically. Specifically, we used the RME Fireface UFX+  [^4] https://rme-audio.de/fireface-ufx.html as the external audio interface and connected it to the ReaInsert. Then, we wrote a ReaScript to automatically send and receive signals from the hardware compressor via the audio interface. To match the level between the DAW and hardware compressor, we normalized all songs to -12 dB and applied a 5 dB input boost and a 5 dB output attenuation. We manually selected 144 widely used parameter combinations for processing after consulting six professional mastering engineers, which are: threshold [-28, -24, -20, -16], attack [0.1, 0.3, 1, 3], release [0.1, 0.4, 0.8, auto], ratio [2, 4, 10]. We additionally recorded 76 other randomly selected combinations as supplementary edge cases. All the audio data was recorded at a sampling rate of 44.1kHz.
Figure 2: The annotation pipeline of Diff-SSL-G-Comp. We utilized various pre-trained models to obtain information on each song’s key, tempo, genre, and instrument.
@@ 3.2 Dataset Statistics
We utilized various pre-trained models to annotate our data, as illustrated in Fig. 2. Specifically, we used the KeyCNN and TempoCNN model  [^5] https://github.com/hendriks73/directional_cnns proposed in [42] to obtain the global music tempo and key information. We split each song into a series of 10s segments and used the Qwen2-Audio [43]  [^6] https://huggingface.co/Qwen/Qwen2-Audio-7B to annotate each segment’s content, which will then be fed to a Llama3 [44]  [^7] https://huggingface.co/meta-llama model to organize and determine the genres and instruments of the whole song.
The statistical results of Diff-SSL-G-Comp on genres, instruments, tempos, and keys are illustrated in Fig. 1 and Fig. 3. From these results, we can conclude that:
  • The majority of genres in our dataset are Rock, Pop, Electronic, and Folk, with a small amount of other uncommon ones like Jazz and Hip Hop.
  • Most used instruments in our dataset are Bass, Drum, Guitar, Vocal, and Synth, with a considerable amount of Piano, String, Keyboard, and Brass. Niche instruments, like Woodwind, are also presented in the dataset.
  • Songs in our dataset are within the range of 70-160 beats per minute (BPM), and the majority of songs are distributed around 110-130 BPM.
  • Most songs in our dataset are in C, D, E, F, G, and A Majors, with a small number of remaining songs evenly distributed across other keys.
@@
Figure 3: Tempo and Key statistics (occurrences) of the unmastered songs used as input signals in Diff-SSL-G-Comp.
@@
Figure 4: Comparison of acoustic and semantic diversities in input signals between Diff-SSL-G-Comp and the existing datasets. The plottings are obtained by applying the PCA algorithm to the SSL representations. We used MERT to extract acoustic embeddings and w2v-BERT 2.0 to extract semantic embeddings. For existing datasets, the compact cluster represents random noises, the diffused cluster represents test signals (sine, square, triangle waves, and their combinations), and the remaining scattered points represent real-world recordings.
@@ 3.3 Dataset Analysis
Unlike existing datasets, which primarily utilize noises and analysis signals, Diff-SSL-G-Comp comprises a collection of diversified real-world unmastered songs as the input signals. To quantify this diversity, we use self-supervised learning (SSL) models to investigate and compare their differences in acoustic and semantic feature spaces, following [38][45], and [46].
Specifically, to analyze the diversity of acoustic features, we leveraged a pre-trained MERT [35]  [^8] https://huggingface.co/m-a-p/MERT-v1-330M model to extract the acoustic representation (the 12th layer is used), which captures various acoustic characteristics such as timbre, style, key, etc. For the semantic diversity analysis, we employed a pre-trained w2v-BERT model [47]  [^9] https://huggingface.co/facebook/w2v-bert-2.0 to generate semantic representations (the last layer is used), capturing melody, lyrics, rhythm, etc. We then applied the Principal Component Analysis (PCA) algorithm to reduce the dimensionality of these representations to two. As illustrated in Fig. 4, most sample points in existing datasets are centered in two distant clusters, where the compact one represents the noise signals, and the diffused one represents the test signals (sine, square, triangle waves, and their combinations), and only a few points scattered aside, representing the real-world instrument recordings. Compared with the existing datasets, Diff-SSL-G-Comp exhibits a broader dispersion in the cluster representing real-world recordings, indicating richer acoustic and semantic characteristic coverage.
@@ 4 Experiments
Table 2: Benchmarking results of existing parametric black-box methods on seen and unseen test sets. The best and second best results of every column are bold and underlined.
@@
Table 3: Benchmarking results of existing grey-box methods on seen and unseen test sets. The best and second best results of every column are bold and underlined.
@@
Table 4: Ablation results of the GCN model trained on different subsets on seen and unseen test sets. The best and second best results of every column in each setting are bold and underlined.
@@
In this section, we conducted benchmark experiments in various black-box and grey-box models and available commercial plugins to verify the effectiveness and facilitate the use of Diff-SSL-G-Comp. We also conducted ablation studies on different data subsets to illustrate the effectiveness of improved data scale and diversity.
@@ 4.1 Experiment Setup
4.1.1 Data Split and Processing
For the train and evaluation data split, we randomly selected 112 songs as the train set and used the remaining 63 songs as the test set. We used our manually selected 144 parameter combinations for training and the seen test distribution. The remaining 76 parameter combinations are used as the unseen test distribution to assess the models’ generalization ability in edge cases.
4.1.2 Training Schedules
All the models are trained using the AdamW [48] optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a initial learning rate of 0.005. The ReduceLROnPlateau Scheduler is used with a factor of 0.5 and a patience of 10000 steps. All the experiments are conducted on a single NVIDIA H200 GPU with a batch size 16 and num workers of 16 for 500K steps. We use the Truncated Backpropagation Through Time (TBPTT) [49] with a 0.01s segment length (4410 samples) to reduce memory costs and enhance training efficiency while maintaining long-term dependencies.
4.1.3 Baselines and Configurations
We use the NablAFx [32] toolbox for conducting benchmarking experiments on baseline systems. Specifically, we use LSTM [2], TCN [3], GCN [4], and S4 [5] for black-box models. The LSTM model is conditioned on direct concatenation (Concat) or time-varying concatenation (TVConcat) [8]. The TCN, GCN, and S4 models are conditioned on FiLM [25], temporal FiLM (TFiLM) [50], tiny temporal FiLM (TTFiLM) [8], and time-varying temporal FiLM (TVFiLM) [8]. We use GreyBoxDRC [7] and two compressor simulation chains proposed in ToneTwist [8] for grey-box models with the original configurations. For commercial plugins, we utilize the available models from Solid State Logic [^10] https://store.solidstatelogic.com/plug-ins/ssl-native-bus-compressor-2, Softube [^11] https://www.softube.com/bus-processor, Overloud [^12] https://www.overloud.com/products/comp-g, and PSPaudioware [^13] https://www.pspaudioware.com/products/psp-busspressor.
4.1.4 Evaluation Metrics
We use the Amphion [51] toolkit for objective evaluation. We use the L1 and Multi-Resolution STFT loss to evaluate the time and frequency-domain errors following ToneTwist [8]. We additionally report the number of trainable parameters to show the model size.

### v2

# Solid State Bus-Comp: A Large-Scale and Diverse Dataset for Dynamic Range Compressor Virtual Analog Modeling

@@ Abstract
Virtual Analog (VA) modeling aims to simulate the behavior of hardware circuits via algorithms to replicate their tone digitally. Dynamic Range Compressor (DRC) is an audio processing module that controls the dynamics of a track by reducing and amplifying the volumes of loud and quiet sounds, which is essential in music production. In recent years, neural-network-based VA modeling has shown great potential in producing high-fidelity models. However, due to the lack of data quantity and diversity, their generalization ability in different parameter settings and input sounds is still limited. To tackle this problem, we present Solid State Bus-Comp, the first large-scale and diverse dataset for modeling the classical VCA compressor — SSL 500 G-Bus. Specifically, we manually collected 175 unmastered songs from the Cambridge Multitrack Library. We recorded the compressed audio in 220 parameter combinations, resulting in an extensive 2528-hour dataset with diverse genres, instruments, tempos, and keys. Moreover, to facilitate the use of our proposed dataset, we conducted benchmark experiments in various open-sourced black-box and grey-box models, as well as white-box plugins. We also conducted ablation studies in different data subsets to illustrate the effectiveness of the improved data diversity and quantity. The dataset and demos are on our project page: https://www.yichenggu.com/SolidStateBusComp/.
@@ 1 Introduction
Virtual Analog (VA) modeling aims to simulate analog audio devices digitally. Dynamic Range Compressor (DRC) is an audio processing module that compresses the dynamics of a track by reducing and amplifying the volumes of loud and quiet sounds, which is essential in music production [1]. VA modeling on DRC is important, but is always considered to be challenging due to its characteristics: non-linear and long temporal dependency.
To model an analog compressor, early DSP-based methods utilized white-box models. Such a model generally comprises a gain computer and a level detector with different algorithm designs [2, 1], which have been well-studied over the years. Apart from this, recent works have also been proposed to explore other potential improvements like increasing computational efficiency [3, 4] and integrating machine-learning techniques for automatic mixing [5, 6]. These developments have led to various achievements in modeling both the entire device [7] and specific components [8, 9].
Although these white-box techniques can deliver high-quality modeling over different devices, the involvement of human experts is often needed, making it hard to automate the modeling process. In recent years, neural-network-based black-box models have developed a lot due to their superior ability to model analog devices in a data-driven way. To be specific,  [10] first proposed an autoencoder model to model various audio effects.  [11] utilized the long short-term memory (LSTM) model for optimizing the long-term dependencies, followed by [12] to further expand into the hyper recurrent neural network (RNN) model with an in-depth comparison between RNN and LSTM models. To utilize the advantages of convolutional neural networks (CNNs),  [13] first employed the WaveNet [14] structure on digital audio effects. Based on this work,  [15] proposed a temporal convolutional network (TCN) with larger receptive fields and huge dilation factors, while  [16] further improved this architecture by integrating the feature-wise linear modulation (FiLM) [17] layers in modeling the parameter conditions. State Space Model (SSM) [18] is another technique to model long-term dependent time series via decomposing a dynamic system into structured state variables.  [19] first employed the S4 blocks in VA modeling, obtaining outstanding performances, followed by  [20] further adopting the latest S6 model [21]. With the development of differentiable digital signal processing (DDSP) [22], works are also proposed to integrate the DSP models’ explainability and efficiency with neural networks. For instance,  [23] proposed differentiable biquad filters for deep learning applications, followed by [24] integrating them with Koopman Networks [25] to operate in a higher-dimensional state space. These advances have also made the neural grey-box models viable. In particular,  [26] utilized the classic white-box DRC [1] design with multilayer perceptrons (MLPs) predicting the parameters in each time frame, followed by  [27] to further simplify the model into parametric Gains for compression and supplementary EQs for non-linear distortion.
Despite the rapid development of VA models, the publicly available datasets are still scarce, with limited data quantity and diversity. Table 1 illustrates the details of the existing datasets regarding DRC. Specifically, early attempts [28] primarily consist of processed short instrument and test signal recordings in a specific parameter setting, tailored for trivial non-parameteric models. SignalTrain [10] first proposed a parametric dataset in modeling the optical compressor LA-2A. It used various randomly generated test signals and a few instrument recordings as the input signals and recorded 20 equally sampled parameter combinations. After that,  [29] proposed the CL-1B dataset with real-world recordings as inputs with more parameter combinations. Recent works like [20] also presented datasets with more diverse devices but often with limited data scale and parameter combinations. Such limitations will significantly constrain the model’s performance, especially when encountering real-world recordings and unseen parameters.
Data scaling has been shown to be effective in many audio-related areas [30, 31, 32, 33]. For instance, Mert [30] utilized a music mixture of 160K hours to scale up a self-supervised representation learning model with 330M parameters, obtaining outstanding performance in music information retrieval; Yue [34] constructed a 650K hours music mixture to train a 7B parameter model for music generation, obtaining state-of-the-art (SOTA) performance; Stable Audio [31] collected 73k hours of audio recordings, leading to SOTA audio generation model with 1B parameters; Emilia [35, 36] presented a 101K hours open-sourced speech dataset, facilitating SOTA speech generations models [32, 37].
@@
Following these previous works, this work presents Solid State Bus-Comp, the first large-scale and diverse dataset for modeling the SSL 500 G-Bus Compressor  [^1] https://solidstatelogic.com/products/stereo-bus-compressor-module. Specifically, we manually selected 175 unmastered real-world songs from the Cambridge Multitrack Library  [^2] https://www.cambridge-mt.com/ms/mtk/ and recorded the compressed signals in 220 parameter combinations, which results in an extensive 2528-hour dataset with diverse genres, instruments, tempos, and keys. To facilitate the use of our dataset, we conducted benchmarking experiments on various open-sourced black-box and grey-box models, as well as available white-box plugins. We also conducted ablation studies on data subsets with different amounts of songs and data scales to illustrate the effectiveness of the improved data diversity and quantity.
@@
Table 1: A comparison of Solid State Bus-Comp with existing VA modeling datasets regarding DRC.
@@
2 Solid State Bus-Comp
@@
Figure 1: Duration statistics (hours) of the unmastered songs used as input signals in Solid State Bus-Comp by genres and instruments.
This section provides the construction details, statistics, and analysis of our proposed Solid State Bus-Comp dataset.
@@ 2.1 Dataset Construction
Solid State Bus-Comp comprises unmastered songs with different genres, instruments, tempos, and keys processed with varying compression parameters. In particular, we manually selected 175 unmastered songs from the Cambridge Multitrack Library \@footnotemark. We used Reaper  [^3] https://www.reaper.fm/ as the Digital Audio Workstation (DAW) to process the data automatically. Specifically, we used the RME Fireface UFX+  [^4] https://rme-audio.de/fireface-ufx.html as the external audio interface and connected it to the ReaInsert. Then, we wrote a ReaScript to automatically send and receive signals from the hardware compressor via the audio interface. To match the level between the DAW and hardware compressor, we normalized all songs to -12 dB and applied a 5 dB input boost and a 5 dB output attenuation. We manually selected 144 widely used parameter combinations for processing after consulting six professional mastering engineers, which are: threshold [-28, -24, -20, -16], attack [0.1, 0.3, 1, 3], release [0.1, 0.4, 0.8, auto], ratio [2, 4, 10]. We additionally recorded 76 other randomly selected combinations as supplementary edge cases. All the audio data was recorded at a sampling rate of 44.1 kHz.
Figure 2: The annotation pipeline of Solid State Bus-Comp. We utilized various pre-trained models to obtain information on each song’s key, tempo, genre, and instrument.
@@ 2.2 Dataset Statistics
We utilized various pre-trained models to annotate our data, as illustrated in Fig. 2. Specifically, we used the KeyCNN  [^5] https://github.com/hendriks73/key-cnn and TempoCNN model  [^6] https://github.com/hendriks73/tempo-cnn proposed in [39]  [^7] https://github.com/hendriks73/directional_cnns to obtain the global music tempo and key information. We split each song into a series of 10s segments and used the Qwen2-Audio [40]  [^8] https://huggingface.co/Qwen/Qwen2-Audio-7B to annotate each segment’s content, which will then be fed to a Llama3 [41]  [^9] https://huggingface.co/meta-llama model to organize the genres and instruments of the whole song.
The statistical results of Solid State Bus-Comp on genres, instruments, tempos, and keys are illustrated in Fig. 1 and Fig. 3. From these results, we can conclude that 1) The majority of genres in our dataset are Rock, Pop, Electronic, and Folk, with a small amount of other uncommon ones like Ambient, Metal, Jazz and Hip Hop; 2) Most used instruments in our dataset are Bass, Drum, Guitar, Vocal, and Synth, with a considerable amount of Piano, String, Keyboard, and Brass. Niche instruments, like Woodwind, are also presented in the dataset; 3) Songs in our dataset are within the range of 70-160 beats per minute (BPM), and the majority of songs are distributed around 110-130 BPM; 4) Most songs in our dataset are in C, D, E, F, G, and A Majors, with a small number of remaining songs evenly distributed across other keys.
@@
Figure 3: Tempo and Key statistics (occurrences) of the unmastered songs used as input signals in our proposed Solid State Bus-Comp. Tempo is in beats per minute (BPM). “M” denotes for “Major” and “m” denotes for “Minor”.
@@
Figure 4: Comparison of acoustic and semantic diversities in input signals between Solid State Bus-Comp and the existing datasets. The plottings are obtained by applying the PCA algorithm to the SSL representations. We used MERT to extract acoustic embeddings and w2v-BERT 2.0 to extract semantic embeddings. For existing datasets, the compact cluster represents random noises, the diffused cluster represents test signals (sine, square, triangle waves, etc), and the remaining scattered points represent real-world recordings.
@@ 2.3 Dataset Analysis
Unlike existing datasets, which primarily utilize noises and analysis signals, Solid State Bus-Comp comprises a collection of diversified real-world unmastered songs as the input signals. To quantify this diversity, we use self-supervised learning (SSL) models to investigate and compare their differences in acoustic and semantic feature spaces, following [35][36], and [33].
Specifically, to analyze the diversity of acoustic features, we leveraged a pre-trained MERT [30]  [^10] https://huggingface.co/m-a-p/MERT-v1-330M model to extract the acoustic representation (the 12th layer is used), which captures various acoustic characteristics such as timbre, style, key, etc. For the semantic diversity analysis, we employed a pre-trained w2v-BERT model [42]  [^11] https://huggingface.co/facebook/w2v-bert-2.0 to generate semantic representations (the last layer is used), capturing melody, lyrics, rhythm, etc. We then applied the Principal Component Analysis (PCA) algorithm to reduce the dimensionality of these representations to two. As illustrated in Fig. 4, most sample points in existing datasets are centered in two distant clusters, where the compact one represents the noise signals, and the diffused one represents the test signals (sine, square, triangle waves, and their combinations), and only a few points scattered aside, representing the real-world instrument recordings. Compared with the existing datasets, Solid State Bus-Comp exhibits a broader dispersion in the cluster representing real-world recordings, indicating richer acoustic and semantic characteristic coverage.
@@ 3 Experiments
Table 2: Benchmarking results of existing parametric black-box methods. The best and second best results are bold and underlined.
@@
Table 3: Benchmarking results of existing grey-box methods. The best and second best results are bold and underlined.
@@
Table 4: Benchmarking results of existing commercial plugins. The best and second best results are bold and underlined.
System L1 ($\downarrow$) M-STFT ($\downarrow$)
Seen Unseen Seen Unseen
Solid State Logic 0.0322 0.0175 0.4489 0.2943
Softube 0.0448 0.0237 0.7069 0.4546
Overloud 0.0326 0.0176 0.4738 0.3253
PSPaudioware 0.0269 0.0145 0.3047 0.2184
Table 5: Ablation results of the GCN model trained on different data subsets. The best and second best results of every column in each setting are bold and underlined.
@@
In this section, we conducted benchmark experiments to verify the effectiveness and facilitate the use of Solid State Bus-Comp. We also conducted ablation studies on different data subsets to illustrate the effectiveness of improved data scale and diversity.
@@ 3.1 Experiment Setup
Data Split and Processing: For the train and evaluation data split, we randomly selected 112 songs as the train set and used the remaining 63 songs as the test set. We used our manually selected 144 parameter combinations for training and the seen test distribution. The remaining 76 parameter combinations are used as the unseen test distribution to assess the generalization ability.
Training Schedules: All the models are trained using the AdamW [43] optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a initial learning rate of 0.005. The ReduceLROnPlateau Scheduler is used with a factor of 0.5 and a patience of 10000 steps. All the experiments are conducted on a single NVIDIA H200 GPU with a batch size 16 and num workers of 16 for 500K steps. We use the Truncated Backpropagation Through Time (TBPTT) [44] with a 0.01s segment length (4410 samples) to reduce memory costs while maintaining long-term dependencies.
Baselines and Configurations: We use the NablAFx [27] toolbox for conducting benchmarking experiments on baseline systems. Specifically, we use LSTM [11], TCN [15], GCN [16], and S4 [19] for black-box models. The LSTM model is conditioned on direct concatenation (Concat) or time-varying concatenation (TVConcat) [38]. The TCN, GCN, and S4 models are conditioned on FiLM [17], temporal FiLM (TFiLM) [45], tiny temporal FiLM (TTFiLM) [38], and time-varying temporal FiLM (TVFiLM) [38]. We use GreyBoxDRC [26] and two compressor simulation chains proposed in ToneTwist [38] for grey-box models with the original configurations. For commercial plugins, we utilize the available models from Solid State Logic [^12] https://store.solidstatelogic.com/plug-ins/ssl-native-bus-compressor-2, Softube [^13] https://www.softube.com/bus-processor, Overloud [^14] https://www.overloud.com/products/comp-g, and PSPaudioware [^15] https://www.pspaudioware.com/products/psp-busspressor. To facilitate reproducible research, all of the modified code and the pre-trained models can be accessed via  [^16] https://drive.google.com/drive/folders/1zf5hnF7XGRW-poo_cqjQthKBeAZx33gd.
Evaluation Metrics: We use the Amphion [46] toolkit for objective evaluation. We use the L1 and Multi-Resolution STFT losses to evaluate the time and frequency-domain errors following ToneTwist [38]. We additionally report the number of trainable parameters to show the model size.